# LLLMs: A Data-Driven Survey of Evolving Research on Limitations of Large Language Models

### v2

@@ Abstract.
Large language model (LLM) research has grown rapidly, along with increasing concern about their limitations such as failures in reasoning, hallucinations, and limited multilingual capability. While prior reviews have addressed these issues, they often focus on individual limitations or consider them within the broader context of evaluating overall model performance. This survey addresses the gap by presenting a data-driven, semi-automated review of research on limitations of LLMs (LLLMs) from 2022 to 2025, using a bottom-up approach. From a corpus of 250,000 ACL and arXiv papers, we extract 14,648 relevant limitation papers using keyword filtering and LLM-based classification, validated against expert labels. Using topic clustering (via two approaches, HDBSCAN+BERTopic and LlooM), we identify between 7 and 15 prominent types of limitations discussed in recent LLM research across the ACL and arXiv datasets. We find that LLM-related research increases nearly sixfold in ACL and nearly fifteenfold in arXiv between 2022 and 2025, while LLLMs research grows even faster, by a factor of over 12 in ACL and nearly 28 in arXiv. Reasoning remains the most studied limitation, followed by generalization, hallucination, bias, and security. The distribution of topics in the ACL dataset stays relatively stable over time, while arXiv shifts toward safety and controllability (with topics like security risks, alignment, hallucinations, knowledge editing), and multimodality between 2022 and 2025. We offer a quantitative view of trends in LLM limitations research and release a dataset of annotated abstracts and a validated methodology, available at: github.com/a-kostikova/LLLMs-Survey.
@@
ccs: Information systems Clustering and classificationccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generationccs: Computing methodologies Information extraction
@@ 1. Introduction
With the explosive growth of large language model (LLM) research and deployment (Wei et al., 2022), questions of the limitations of LLMs (LLLMs) have also gathered increased interest, ranging from reasoning failures (Park et al., 2024), social bias (Lin et al., 2022), hallucinations (Xu et al., 2024a), difficulty in handling long contexts (Li et al., 2023b), and many more. Understanding where LLMs fail is essential for knowing how and whether they can be safely and effectively used in real-world settings, especially as LLMs are increasingly deployed in safety-sensitive domains such as healthcare, education, finance, and law (Chen et al., 2024a). Moreover, tracking how these failure modes evolve over time helps reveal whether the fast-paced research landscape is addressing them, overlooking them, or exposing new ones, offering a clearer picture of where further research is most needed.
However, given the sheer size of LLM research, with thousands of published research papers every year (even when limited to highly rated outlets), it is challenging to maintain an up-to-date overview of LLLMs research using traditional, manual literature review techniques. Accordingly, prior reviews on LLLMs mostly focus on specific limitations, such as reasoning (Li et al., 2025; Xu et al., 2025), or examine limitations within the broader context of evaluating overall model capabilities (Srivastava et al., 2022; Chang et al., 2024). To date, the field still misses an overview that covers the more recent LLM research between 2022 and now and cuts across limitations. Our review is an attempt to provide this high-level overview.
@@
We observe four main results. i) LLLMs research has grown rapidly, outpacing even the growth of LLM research overall. The number of LLM-related papers has grown by a factor of nearly 6 in ACL and nearly 15 in arXiv between 2022-2025, reaching almost 80% of all crawled ACL papers and roughly 30% of all crawled arXiv papers; LLLMs papers have increased even more sharply, by a factor of over 12 in ACL and 28 in arXiv, accounting for more than 30% of LLM papers in Q1 of 2025. ii) Within LLLMs research, reasoning limitations are the most prominent, with generalization, hallucination, bias, and security as further important concerns. iii) The distribution of limitations appears relatively stable in the ACL dataset, whereas the arXiv dataset shows a rise in concern for topics related to safety and controllability (e.g., Security Risks, Alignment Limitations, Knowledge Editing, Hallucination) as well as Multimodality. (iv) Despite substantial methodological differences between HDBSCAN and LlooM, we observe topical overlap in several of the biggest clusters (e.g., Reasoning, Hallucination, Security Risks) across both approaches, with broadly similar trend patterns, suggesting that these findings are reliable.
@@ 2.1. Surveys of Large Language Models
In parallel to cross-domain surveys, a number of studies have investigated how LLMs are being adopted and evaluated in specific fields. In the medical domain, surveys examine the effectiveness of LLMs in clinical summarization and diagnostic reasoning (Kim et al., 2025; Van Veen et al., 2024), as well as challenges related to hallucination and factual consistency in medical question answering (Pham and Vo, 2024). Other works explore the capabilities and limitations of LLMs in capturing cultural commonsense knowledge (Shen et al., 2024) and scientific research processes (Luo et al., 2025; Eger et al., 2025). In recommendation systems, LLMs have been studied as both retrieval and generation engines (Wu et al., 2024), while in information retrieval, surveys highlight their use in query expansion, passage ranking, and answer synthesis (Zhu et al., 2023). LLMs have also been applied to software engineering tasks such as code generation and bug fixing, with systematic reviews discussing both their potential and practical limitations (Hou et al., 2024). Beyond text-based applications, LLMs have been integrated with structured resources such as knowledge graphs (Pan et al., 2024; Jin et al., 2024) and studied in the context of autonomous agent design (Wang et al., 2024c).
@@ 2.2. Surveys on Limitations of LLMs
As LLMs are increasingly deployed in real-world applications, a growing body of research has emerged to examine their limitations from different capability-oriented perspectives. One prominent area of concern are hallucinations, where recent surveys investigate underlying causes and mitigation strategies in both text generation (Ji et al., 2023; Huang et al., 2025; Tonmoy et al., 2024) and multimodal contexts (Liu et al., 2024; Sahoo et al., 2024, [n. d.]). Another major focus is reasoning, with surveys analyzing the development of novel techniques(Plaat et al., 2024; Huang and Chang, 2022) such as chain-of-thought prompting (Chen et al., 2025), reinforced reasoning (Xu et al., 2025; Li et al., 2025), and mathematical problem-solving (Ahn et al., 2024). In parallel, the trustworthiness and reliability of LLMs have been studied through the lens of fairness, transparency, and calibration (Sun et al., 2024), while other works concentrate on security and privacy threats including adversarial vulnerabilities and data leakage (Yao et al., 2024; Das et al., 2025). A related thread explores ethical risks in LLM-based agents, particularly concerning safety, misuse, and human interaction (Jiao et al., 2024; Kumar et al., 2024; Gan et al., 2024).
@@ 2.3. LLMs as Analytical Tools for Scientific Literature
We apply a partially automated pipeline, relying on LLMs to filter the papers included in our survey and providing embeddings for clustering. Such methods have to be applied with care to avoid being misled because of the very limitations this survey is supposed to study. In developing our methodology, we rely on a growing literature of LLMs being used as instruments for analyzing scientific literature (Eger et al., 2025). Several recent approaches employ LLMs for topic modeling, semantic clustering, and concept induction, enabling more interpretable organization of large-scale corpora (Zhang et al., 2023b; Diaz-Rodriguez, 2025; Lam et al., 2024; Feng et al., 2024; Pattnaik et al., 2024). This has led to a growing number of systematic reviews that use BERTopic and related methods to map research landscapes across areas such as generative AI, information assurance, LLM applications, and research impact evaluation (Gana et al., 2024; Ding et al., 2024; Gupta et al., 2024; Arsalan et al., 2025; Leiva-Araos et al., 2025).
In addition to these analytical techniques, other efforts focus on automating synthesis of results across papers. Systems such as SurveyX and AutoSurvey generate draft surveys from large paper collections (Liang et al., 2025; Wang et al., 2024b), while tools like LitLLM (Agarwal et al., 2024) and PaSa (He et al., 2025) support LLM-based retrieval, summarization, and exploration of academic texts.
@@ 3. Methodology
Fig. 1 illustrates the method for our systematic literature review. We begin by retrieving papers from arXiv and ACL (Section 3.1), filter according to keywords (Section 3.2), filter papers further by classifying their abstracts with an LLM (Section 3.3), and finally cluster the papers (Section 3.5). At each step of the analysis, we perform validations to ensure the robustness of our results: the keyword list is obtained with an iterative refinement procedure, the LLM classification step is validated against a gold standard of 445 human-annotated papers, and we use two distinct clustering methods for comparison (HDBSCAN+BERTopic and LlooM). In the remainder of this section, we describe each step in more detail.
@@ 3.1. Data Retrieval
Our initial dataset includes all academic papers published between January 2022 and March 2025, sourced from the ACL Anthology and arXiv. ArXiv captures (potentially) non–peer-reviewed research that closely tracks current developments, while ACL venues reflect peer-reviewed work and remain the primary publication outlets for NLP research, where much of the foundational work on LLMs originated. The time frame was chosen to capture the year preceding the release of ChatGPT as well as all subsequent research on LLMs (Zhao et al., 2023).
For ACL Anthology, we scrape conference pages for AACL 2022–2023, ACL 2022–2024, EACL 2023–2024, EMNLP 2022–2024, ICLR 2022–2024, NAACL 2022 and 2024, and TACL 2022–2024 as the premier NLP venues. [^2] We exclude other tracks and venues such as workshops, system demonstrations, tutorials, shared tasks, and task-specific venues (e.g., SemEval, CoNLL, WMT) to maintain a focus on high-impact research from general-purpose NLP conferences. For arXiv, we retrieve papers from the categories of Computation & Language (cs.CL), Machine Learning (cs.LG), Artificial Intelligence (cs.AI), and Computer Vision (cs.CV) because these communities are closest to LLM research. However, we note that many papers are classified by the authors into multiple arXiv categories, so that research areas beyond these initial ones are covered as well (see Section 4.4.3 for more details). Each entry includes metadata such as title, publication date, author information, download link, and abstract, with arXiv papers also containing all assigned categories. We use titles and abstracts for keyword filtering and clustering, as they capture a paper’s main claims and contributions and are well-suited for large-scale automated analysis. The final crawled dataset includes 245,835 papers (18,578 papers for ACL, 227,257 for arXiv). Fig. 2 shows raw numbers of crawled ACL and arXiv papers over time.
@@ 3.2. Keyword-Based Filtering
    We use TNT-KID (Martinc et al., 2022) to generate initial keywords for each abstract.
@@
This results in a list of 90 keywords (19 unigrams, 44 bigrams, 16 trigrams, and 11 four-grams). Overall, the final keyword set covers key aspects of LLM research (see the full list in Section A supplementary material):
@@
We keep only those papers that contain at least one of the keywords in their title or abstract. This results in 64,110 papers (8,635 for ACL, 55,475 for arXiv). A breakdown of crawled and filtered papers across sources for each year is provided in Table 1. While this filtering step does not strictly isolate LLM-focused papers, we observe that the proportion of papers passing the filter increases over time, which may reflect a growing focus on LLMs in the broader NLP research landscape. We examine this pattern in more detail in 4.2, where we analyze trends in both LLM and LLLMs papers after additional filtering.
@@ 3.3.1. Human Annotation Task
For papers rated 2-5, annotators highlighted textual evidence pointing to the limitation and its domain, hereafter referred to as ?evidence?. See Table 13 in the supplementary material for representative examples of annotated papers and highlighted evidence.
@@
The annotation process included multiple rounds, involving two professors (natural language processing and machine learning), one PhD student (NLP), and one Master’s student (computer science). Initial annotations showed moderate inter-annotator agreement (0.27 standard Cohen’s Kappa, and 0.62 weighted), which was improved by further rounds (0.57, and 0.75, respectively), indicating substantial agreement in the final version. Overall, the annotation process covered 445 samples (where 195 were annotated solely by a Master’s student after all discussion rounds).
@@
We compare the evidence highlighted based on a representation as BIO-tagged sequences, where tokens are labeled as B-EVID, I-EVID, or O (beginning, inside or outside evidence, respectively). In our agreement analysis, we include any paper that is annotated by at least two annotators, which was the case for 250 of the 455 annotated papers. We compute agreement by means of the macro-averaged F1 score across annotator pairs, excluding cases where neither annotator selected evidence.

### v3

@@ Abstract.
Large language model (LLM) research has grown rapidly, along with increasing concern about their limitations. In this survey, we conduct a data-driven, semi-automated review of research on limitations of LLMs (LLLMs) from 2022 to early 2025 using a bottom-up approach. From a corpus of 250,000 ACL and arXiv papers, we identify 14,648 relevant papers using keyword filtering, LLM-based classification, validated against expert labels, and topic clustering (via two approaches, HDBSCAN+BERTopic and LlooM). We find that the share of LLM-related papers increases over fivefold in ACL and nearly eightfold in arXiv between 2022 and 2025. Since 2022, LLLMs research grows even faster, reaching over 30% of LLM papers by 2025. Reasoning remains the most studied limitation, followed by generalization, hallucination, bias, and security. The distribution of topics in the ACL dataset stays relatively stable over time, while arXiv shifts toward security risks, alignment, hallucinations, knowledge editing, and multimodality. We offer a quantitative view of trends in LLLMs research and release a dataset of annotated abstracts and a validated methodology, available at: github.com/a-kostikova/LLLMs-Survey.
@@
$c$s: Information systems Clustering and classificationccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generationccs: Computing methodologies Information extraction
@@ 1. Introduction
With the explosive growth of large language model (LLM) research and deployment (Wei et al., 2022), questions of the limitations of LLMs (LLLMs) have also gathered increased interest, ranging from reasoning failures (Park et al., 2024), social bias (Lin et al., 2022), hallucinations (Xu et al., 2024), difficulty in handling long contexts (Li et al., 2023b), and many more. Understanding where LLMs fail is essential for knowing how and whether they can be safely and effectively used in real-world settings, especially as LLMs are increasingly deployed in safety-sensitive domains such as healthcare, education, finance, and law (Chen et al., 2024). Moreover, tracking how these failure modes evolve over time helps reveal whether the fast-paced research landscape is addressing them, overlooking them, or exposing new ones, offering a clearer picture of where further research is most needed.
However, given the sheer size of LLM research, with thousands of published research papers every year (even when limited to highly rated outlets), it is challenging to maintain an up-to-date overview of LLLMs research using traditional, manual literature review techniques. Accordingly, prior reviews on LLLMs mostly focus on specific limitations, such as reasoning (Li et al., 2025; Xu et al., 2025a), or examine limitations within the broader context of evaluating overall model capabilities (Srivastava et al., 2022; Chang et al., 2024). To date, the field still misses an overview that covers the more recent LLM research between 2022 and now and cuts across limitations. Our review is an attempt to provide this high-level overview.
@@
We observe four main results. i) LLLMs research has grown rapidly, outpacing even the growth of LLM research overall. The share of LLM-related papers has grown by a factor of over five in ACL and nearly eight in arXiv between 2022-2025, reaching almost 80% of all crawled ACL papers and roughly 30% of all crawled arXiv papers; LLLMs papers have increased even more sharply, by a factor of over 12 in ACL and 28 in arXiv, accounting for more than 30% of LLM papers in Q1 of 2025. ii) Within LLLMs research, reasoning limitations are the most prominent, with generalization, hallucination, bias, and security as further important concerns. iii) The distribution of limitations appears relatively stable in the ACL dataset, whereas the arXiv dataset shows a rise in concern for topics related to safety and controllability (e.g., Security Risks, Alignment Limitations, Knowledge Editing, Hallucination) as well as Multimodality. (iv) Despite substantial methodological differences between HDBSCAN and LlooM, we observe topical overlap in several of the biggest clusters (e.g., Reasoning, Hallucination, Security Risks) across both approaches, with broadly similar trend patterns, suggesting that these findings are reliable.
@@ 2.1. Surveys of Large Language Models
In parallel to cross-domain surveys, a number of studies have investigated how LLMs are being adopted and evaluated in specific fields. In the medical domain, surveys examine the effectiveness of LLMs in clinical summarization and diagnostic reasoning (Kim et al., 2025; Van Veen et al., 2024), as well as challenges related to hallucination and factual consistency in medical question answering (Pham and Vo, 2024). Other works explore the capabilities and limitations of LLMs in capturing cultural commonsense knowledge (Shen et al., 2024) and scientific research processes (Luo et al., 2025; Eger et al., 2025). In recommendation systems, LLMs have been studied as both retrieval and generation engines (Wu et al., 2024), while in information retrieval, surveys highlight their use in query expansion, passage ranking, and answer synthesis (Zhu et al., 2023). LLMs have also been applied to software engineering tasks such as code generation and bug fixing, with systematic reviews discussing both their potential and practical limitations (Hou et al., 2024). Beyond text-based applications, LLMs have been integrated with structured resources such as knowledge graphs (Pan et al., 2024; Jin et al., 2024) and studied in the context of autonomous agent design (Wang et al., 2024a).
@@ 2.2. Surveys on Limitations of LLMs
As LLMs are increasingly deployed in real-world applications, a growing body of research has emerged to examine their limitations from different capability-oriented perspectives. One prominent area of concern are hallucinations, where recent surveys investigate underlying causes and mitigation strategies in both text generation (Ji et al., 2023; Huang et al., 2025; Tonmoy et al., 2024) and multimodal contexts (Liu et al., 2024; Sahoo et al., 2024, ). Another major focus is reasoning, with surveys analyzing the development of novel techniques(Plaat et al., 2024; Huang and Chang, 2022) such as chain-of-thought prompting (Chen et al., 2025), reinforced reasoning (Xu et al., 2025a; Li et al., 2025), and mathematical problem-solving (Ahn et al., 2024). In parallel, the trustworthiness and reliability of LLMs have been studied through the lens of fairness, transparency, and calibration (Sun et al., 2024), while other works concentrate on security and privacy threats including adversarial vulnerabilities and data leakage (Yao et al., 2024; Das et al., 2025). A related thread explores ethical risks in LLM-based agents, particularly concerning safety, misuse, and human interaction (Jiao et al., 2024; Kumar et al., 2024; Gan et al., 2024).
@@ 2.3. LLMs as Analytical Tools for Scientific Literature
We apply a partially automated pipeline, relying on LLMs to filter the papers included in our survey and providing embeddings for clustering. Such methods have to be applied with care to avoid being misled because of the very limitations this survey is supposed to study. In developing our methodology, we rely on a growing literature of LLMs being used as instruments for analyzing scientific literature (Eger et al., 2025). Several recent approaches employ LLMs for topic modeling, semantic clustering, and concept induction, enabling more interpretable organization of large-scale corpora (Zhang et al., 2023c; Diaz-Rodriguez, 2025; Lam et al., 2024; Feng et al., 2024; Pattnaik et al., 2024). This has led to a growing number of systematic reviews that use BERTopic and related methods to map research landscapes across areas such as generative AI, information assurance, LLM applications, and research impact evaluation (Gana et al., 2024; Ding,Qinxu et al., 2024; Gupta et al., 2024; Arsalan et al., 2025).
In addition to these analytical techniques, other efforts focus on automating synthesis of results across papers. Systems such as SurveyX and AutoSurvey generate draft surveys from large paper collections (Liang et al., 2025; Wang et al., 2024c), while tools like LitLLM (Agarwal et al., 2024) and PaSa (He et al., 2025) support LLM-based retrieval, summarization, and exploration of academic texts.
Closer to our work, recent studies use automated methods to analyze research limitations at the paper or corpus level, primarily in support of peer review. Al Azher et al. (2025) introduce datasets and methods for extracting and generating study limitations, while Azher et al. (2024) apply LLM-assisted topic modeling to limitation sections. Xu et al. (2025b) evaluate whether LLMs can identify critical limitations in AI papers. In contrast, our work studies LLLMs themselves as the research object.
@@ 3. Methodology
Fig. 1 illustrates the method for our systematic literature review. We begin by retrieving papers from arXiv and ACL (Section 3.1), filter according to keywords (Section 3.2), filter papers further by classifying their abstracts with an LLM (Section 3.3), and finally cluster the papers (Section 3.5). At each step of the analysis, we perform validations to ensure the robustness of our results: the keyword list is obtained with an iterative refinement procedure, the LLM classification step is validated against a gold standard of 445 human-annotated papers, and we use two distinct clustering methods for comparison (HDBSCAN+BERTopic and LlooM), complemented by stratified topic-level human evaluation of cluster assignments. In the remainder of this section, we describe each step in more detail.
@@ 3.1. Data Retrieval
Our initial dataset includes academic papers published between January 2022 and March 2025, sourced from arXiv as a broad, multi-field research corpus, and from the ACL Anthology as a narrower, NLP-focused set. ArXiv captures preprint research that closely tracks current developments and often appears months before formal publication, making it well suited for analyzing fast-moving research trends in LLMs, while ACL venues reflect peer-reviewed work and remain the primary publication outlets for NLP research, where much of the foundational work on LLMs originated and where consistent multi-year coverage enables longitudinal analysis. [^2] In the arXiv dataset, some papers later appear in ACL venues; we retain them to preserve arXiv’s role as a source of fast-moving, venue-agnostic research. The time frame was chosen to capture the year preceding the release of ChatGPT as well as all subsequent research on LLMs (Zhao et al., 2023).
For ACL Anthology, we scrape conference pages for AACL 2022–2023, ACL 2022–2024, EACL 2023–2024, EMNLP 2022–2024, ICLR 2022–2024333Although ICLR is not an ACL venue nor part of the ACL Anthology, we include it in this group to represent a major peer-reviewed machine learning conference with substantial LLM research, and retain the ?ACL? label for brevity., NAACL 2022 and 2024, and TACL 2022–2024 as the premier NLP venues. [^4] We exclude other tracks and venues such as workshops, system demonstrations, tutorials, shared tasks, and task-specific venues (e.g., SemEval, CoNLL, WMT) to maintain a focus on high-impact research from general-purpose NLP conferences. For arXiv, we retrieve papers from the categories of Computation & Language (cs.CL), Machine Learning (cs.LG), Artificial Intelligence (cs.AI), and Computer Vision (cs.CV) because these communities are closest to LLM research. However, we note that many papers are classified by the authors into multiple arXiv categories, so that research areas beyond these initial ones are covered as well (see Section 4.3.3 for more details). Each entry includes metadata such as title, publication date, author information, download link, and abstract, with arXiv papers also containing all assigned categories. We use titles and abstracts for keyword filtering and clustering, as they capture a paper’s main claims and contributions and are consistently available across venues, making them well suited for large-scale automated analysis. Including additional sections (e.g., introductions or related work) could introduce noise, increase preprocessing complexity due to heterogeneous document structures, and raise the cost of large-scale LLM-based annotation.
The final crawled dataset includes 245,835 papers (18,578 papers for ACL, 227,257 for arXiv). Fig. 2 shows raw numbers of crawled ACL and arXiv papers over time.
@@ 3.2. Keyword-Based Filtering
    We use TNT-KID (Martinc et al., 2022) to generate initial keywords for each abstract. TNT-KID is a transformer-based keyword tagger that follows a two-stage training process: language model pretraining on large unlabeled text, followed by fine-tuning on labeled keyword data. As reported by Giarelis and Karacapilidis (2024), TNT-KID is among the top-performing keyphrase extraction methods and provides ready-to-use pretrained models. We use the off-the-shelf TNT-KID model fine-tuned on the KP20k dataset of scientific abstracts and titles.
@@
This results in a list of 90 keywords (19 unigrams, 44 bigrams, 16 trigrams, and 11 four-grams). Overall, the final keyword set covers key aspects of LLM research (see the full list in Section A of the online supplementary material):
@@
We keep only those papers that contain at least one of the keywords in their title or abstract. This results in 64,110 papers (8,635 for ACL, 55,475 for arXiv). A breakdown of crawled and filtered papers across sources for each year is provided in Table 1. While this filtering step does not strictly isolate LLM-focused papers, it serves as an initial inclusion step; its output is refined by subsequent LLM-based classification. Even at this stage, we observe that the proportion of papers passing the filter increases over time, which may reflect a growing focus on LLMs in the broader NLP research landscape. We examine this pattern in more detail in 4.2, where we analyze trends in both LLM and LLLMs papers after additional filtering.
@@ 3.3.1. Human Annotation Task
For papers rated 2–5, annotators highlighted textual evidence pointing to the limitation and, where explicitly expressed in the text, its general type (i.e., the kind of limitation discussed), hereafter referred to as ?evidence?. See Table 2 in the online supplementary material for representative examples of annotated papers and highlighted evidence.
@@
In our agreement analysis, we include all papers annotated by at least two annotators (up to four). Agreement is computed pairwise for each annotator pair over the subset of papers they both annotated and averaged across all pairs.
The annotation process included three rounds, involving two professors (natural language processing and machine learning), one PhD student (NLP), and one Master’s student (computer science), all of whom are fluent English speakers with at least C1 proficiency. Disagreements were addressed through discussion and refinement of the annotation guidelines between rounds, without per-instance adjudication. Initial annotations showed moderate inter-annotator agreement (0.27 standard Cohen’s Kappa, and 0.62 weighted), which was improved by further rounds (0.57, and 0.75, respectively), indicating substantial agreement in the final version. Overall, the annotation process covered 445 samples (where 195 were annotated solely by a Master’s student after all discussion rounds).
@@
We compare the highlighted evidence using BIO-tagged sequences, where tokens are labeled as B-EVID, I-EVID, or O (beginning, inside or outside evidence, respectively). For each paper annotated by at least two annotators (250 of 455 abstracts), we compute pairwise inter-annotator F1 scores at the token level. For each annotator pair and paper, we compare BIO-tagged token sequences, treating B-EVID and I-EVID as positive labels and O as the negative label, and compute F1, excluding annotator pairs for which neither annotator selected any evidence. We report macro-average F1 across all valid annotator pairs.